Ive totally lost myself with this, so if there is anyone that can help please

2log[4](x)+log[4](3)=log[4](x)-log[4](2)

All the 4s are log base 4
Thank you

Answer by stanbon(57940)

2log[4](x)+log[4](3)=log[4](x)-log[4](2)
-----
log4(x^2) + log4(3) = log4[x/2]
---
log4(3x^2) = log4(x/2)
---
3x^2 = x/2
6x^2 = x
6x^2-x = 0
x(6x-1) = 0
---
x cannot be zero, so x = 1/6 is the only answer.
---
Let me know if you do not understand all this.
Cheers,
Stan H.
=============